I am trying to help me 12 year old with his word problems, but I think I am going in the wrong direction and will confuse him. The question is " Nine less eight times a number is one less than negative two times the difference of three times the number and six.
8n-9=3n + 6 - 1-2n
Please help --- am i on the right track?? Thanks so much!

Answer by stanbon(57361)

Nine less eight times a number is one less than negative two times the difference of three times the number and six.
-------
Equation:
Let the number be "x":
9 - 8x = -2(3x-6)-1
-----
9 - 8x = -6x+12-1
---
9 - 8x = -6x+11
2x = -2
x = -1
================
cheers,
Stan H.